# The Dark Side of Rich Rewards: Understanding and Mitigating Noise in VLM Rewards

### v4

@@ A.2.1 Convergence time on a sparse-reward landscape
Proof: let $\vec{X_{1}}$, $\vec{X_{2}}$, … $\vec{X_{T}}$ be IID random unit vectors uniformly distributed on a $(d-1)$-dimensional sphere $\mathcal{S}^{d-1}\subset\mathbb{R}^{d}$, where $\vec{X_{i}}=(X_{i1},...,X_{id})$ and $|\vec{X_{i}}|^{2}=\sum_{j=1}^{d}X^{2}_{ij}=1$. Let $\vec{S_{T}}:=\sum_{i=1}^{T}\vec{X_{i}}$. By a $n$-dimensional cosine rule we have
@@
i.e. $\mathbb{E}[|\vec{S}_{T}|]\sim\sqrt{T}$. When a policy is randomly initialized with $\theta_{0}$, the distance $D$ to a goal policy, $D:=||\theta_{goal}-\theta_{0}||$ is fixed and is the distance the random walk needs to travel ($\mathbb{E}[|\vec{S}_{T}|]=D\sim\sqrt{T_{D}}$), so the characteristic time needed to travel this distance $T_{D}\sim D^{2}$ as we have shown above.
@@
See 4.2 Proof: the expected time taken for each of the sub-tasks then scales with ${}d_{i}^{2}$ respectively (Lemma A.2), and we have
@@
because $d_{i}>0\forall i$. We can also work out that the upper bound for this improvement is a factor of $n-1$ by invoking the Cauchy–Schwarz inequality $\left(\sum_{i=1}^{n}u_{i}^{2}\right)\left(\sum_{i=1}^{n}v_{i}^{2}\right)\geq\left(\sum_{i=1}^{n}u_{i}v_{i}\right)^{2}$:
@@
and the equality sign holds (indicating maximal improvement) when $d_{1}=d_{2}=...=d_{n-1}$. The intuitive interpretation is that the divide-and-conquer approach is the most effective when the task is divided evenly into subtasks. Additionally, the presence of $p(l_{k})$ introduces some lexical-level tolerance in the learning process and results in a considerably larger $\delta$-close radius which could result in further improvement in the convergence rate. However, as we will discuss in the next section, introducing too much tolerance could harm the convergence rate rather than improve it.
A.2.2 False Positive and the Violation of Pessimistic Property of Heuristics
See 4.7
@@ A.2.2 False Positive and the Violation of Pessimistic Property of Heuristics
    Express the Bellman backup for $h$: $(\mathcal{B}h)(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)h(s^{\prime})$
@@
See 4.8 We analyze the convergence speed of policy learning by examining the performance gap, which is defined as the difference between the optimal value of the initial state $s_{0}$, $V^{*}(s_{0})$, and the value of the initial state under an arbitrary policy $\pi$, $V^{\pi}(s_{0})$. Specifically, we focus on deriving an upper bound for this performance gap. The key intuition is that a smaller upper bound implies faster convergence to the optimal policy, as fewer iterations of policy updates will be required to reach the optimum.
@@
Theorem A.3 (Performance Gap Decomposition (Cheng et al., 2021)).
@@
We will not provide a proof for Theorem A.3 in this paper. Please refer to (Cheng et al., 2021) for details. The theorem demonstrates that the performance gap can be elegantly decomposed into two components: a regret term and a bias term such that:
@@
Cheng et al. (2021) have further derived the upper bound for the bias term, which we present here as a lemma. We omit the proof for brevity; for detailed proof, please refer to (Cheng et al., 2021).
Lemma A.4 (Upper Bound of the Bias Term).
@@
    To begin with, we need to use another lemma from Cheng et al. (2021), shown as follows:
    Lemma A.6 (Bellman backup of reshaped MDP (Cheng et al., 2021)).
@@
    For brevity, we do not include the proof for Lemma A.6.
@@
    Then, according to Lemma A.6, we can get
@@
    Let $\pi$ denote the greedy policy of $\arg\max_{a}(\mathcal{B}h)(s,a)$ and then trace back to Equation 8, we have (~$(\widetilde{\mathcal{B}}h)(s,a)=(\mathcal{B}h)(s,a)$, that means
@@
    $\geq h(s_{0})\quad\text{[direct result of $h$ being pessimistic]}$] (39)
    Generalization to Any State: The lemma from Cheng et al. (2021) (Lemma A.6) is stated in terms of any starting state $s_{0}$. Therefore, we can replace $s_{0}$ with any state $s\in\mathcal{S}$ in our analysis. Thus we get $\forall s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
    Lemma A.5 implies that when $h$ is pessimistic, $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\leq 0$ because $\widetilde{V}^{*}(s)\geq h(s)$ for all $s$. Therefore, the bias error is only bounded by $B_{1}$ when $h$ is pessimistic, i.e., $\mathrm{Bias}(h,\beta,\pi)\leq B_{1}$.
@@
    We show briefly that, unlike the underestimation error, the overestimation error does not have a closed-form upper bound expression. The difficulty, as pointed out by (Cheng et al., 2021), originates from the trajectory-dependence on $\mathbb{E}_{\rho^{\pi}}[\cdot]$ as the $l_{\infty}$ approximation error here can be difficult to control in large state spaces. It is possible that, upon picking up falsely high $h$ states, $\rho^{\pi}$ gets further distorted away from $\rho^{\pi^{*}}$ and accumulates more falsely high $h$ states. In other words, this trajectory dependence makes $B_{2}$ prone to a feedback loop of accumulating overestimation errors and resulting in a much larger upper bound for the bias.
A.3 Additional Notes on Theoretical Setting
A.3.1 Convergence Guarantee of HuRL Compared With Potential-Based Reward Shaping
Previous work on potential-based reward shaping (PBRS) has established that learning a policy on a reshaped MDP can converge to the optimum of the original MDP, as proved by Ng et al. (1999). They proved that the optimal $Q$ value of the reshaped MDP is equivalent to the optimal $Q$ value of the original MDP minus a state-dependent function. Consequently, for an optimal policy defined as $\pi^{*}=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a)$, this additional state-dependent value does not alter the policy. Thus, the optimal policy of the original MDP $\mathcal{M}$ is also the optimal policy for the reshaped $\widetilde{\mathcal{M}}$. However, the theoretical convergence for the reshaped MDP within the framework of Heuristic-Guided Reinforcement Learning (HuRL) has not been proven with the same rigor.
The authors of HuRL made a trick which involves manipulating the coefficient $\beta$, which scales the heuristic-based auxiliary rewards. If $\beta$ increases gradually from 0 to 1 over the training process, the agent effectively transitions from interacting with the modified MDP back to the original MDP. This method ensures that, at the end of training, the agent is exactly solving the original problem. Moreover, according to the Blackwell optimal property (Chen et al., 2018), convergence can occur before $\beta$ reaches 1, thereby allowing the trained policy to maintain optimality in the original MDP under HuRL.
@@ A.3.1 Convergence Guarantee of HuRL Compared With Potential-Based Reward Shaping
A.4 Implementation Details of the Experiments
A.4.1 Environment Details
We describe each testing environment used in our experiments. More details introduction can be found in on the official project homepage of each benchmark (Chevalier-Boisvert et al., 2018; Bellemare et al., 2013; Hafner, 2021).
A.4.2 Instruction-Guide Procedure Details
@@
The VLM-based reward model will have a pointer to the sequence of the instruction sentence, starting at the first sentence. For original models Pixl2R and ELLM-, we follow the setting in there original work where for each instruction sentence (yes, the full instruction essay will be split into multiple sentences and treat each sentence as atomic instruction $l_{k}$), the reward model will have a maximum cap of rewards (2.0) it can assign to the agent in one episode. When the cap is reached, the reward model will move its pointer to the next instruction sentence. For the BiMI reward model, the reward model will move its pointer to the next instruction sentence when the binary signal is triggered. Below is the pseudo-code for the instruction-following RL training procedure for both Pixl2R and ELLM- models.
Algorithm 1 ELLM and Pixl2R pseudo-code for instruction-following RL
1:Initialize policy network $\pi_{\theta}$
2:Initialize value network $V_{\phi}$
3:Setup VLM-based reward model $E(\cdot)$
4:Split instruction essay into sentences $\{l_{1},l_{2},...,l_{K}\}$
5:Initialize instruction pointer $p=1$
6:Initialize cumulative VLM reward $r_{\text{cum}}=0$
7:Initialize cumulative VLM reward threshold $q$
8:Initialize replay buffer $\mathcal{D}$
9:Initialize agent trajectory memory queue $\tau$ with length $W$Foreach episode
[…]
11:Select action $a_{t}\sim\pi_{\theta}(a_{t}|s_{t})$
12:Execute $a_{t}$, observe next state $s_{t+1}$ and extrinsic reward $r^{e}_{t}$
13:Enqueue $(s_{t},a_{t},r_{t},s_{t+1})$ in $\tau$
14:Compute VLM reward:
15:$r^{v}_{t}=\frac{E(\tau)\cdot E(l_{p})}{\|E(\tau)\|\|E(l_{p})\|}$triangleright Apply BiMI here if needed
16:Combine rewards: $r_{t}=r^{e}_{t}+(1-\beta)\gamma r^{v}_{t}$triangleright $\beta$ is a scaling factor
17:Store $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{D}$
[…]
19:$p\leftarrow\min(p+1,K)$
[…]
21:Sample mini-batch $\{(s_{j},a_{j},r_{j},s_{j+1})\}$ from $\mathcal{D}$
22:Compute TD errors:
23:$\delta_{j}=r_{j}+\gamma V_{\phi}(s_{j+1})-V_{\phi}(s_{j})$
24:Update value network:
25:$\phi\leftarrow\phi+\alpha_{v}\sum_{j}\delta_{j}\nabla_{\phi}V_{\phi}(s_{j})$
26:Update policy network:
[…]
A.4.3 Finetuning VLM-based Reward Models
@@
In contrast to previous work on instructing following RL where they rely on hand-crafted oracle multimodal reward models, we use actual pretrained VLMs to generate reward signals. 2 VLM backbone models are used in our experiments: 1) CLIP (Radford et al., 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al., 2022), pretrained by video-text pairs. In particular, Pixl2R uses CLIP because it only uses the single latest frame as input. In contrast, ELLM- takes a slice of trajectory (i.e., multiple frames) as input, and thus uses either X-CLIP or CLIP with additional RNN encoder as the reward model.
@@
Due to the cartoonish and abstract visuals of the testing environments, we need to further fine-tune the VLMs to adapt to this new visual domain. We use well-trained expert agents based on Moon et al. (2023) to generate expert trajectories for the Crafter environments and annotate them with instructions using internal information from the game engine. For Minigrid environments, we use classical search-based planning robots to generate expert trajectories and annotate them with the corresponding task instructions. For Montezuma’s Revenge, we manually annotate the expert trajectories.
For Minigrid and Crafter, we have 80,000 training pairs, while for Montezuma’s Revenge, we have around 300 training pairs. These training data are of high quality, as we have made every effort to avoid false positive rewards due to poor training data quality. To enhance our models’ robustness, we also employed contrastive learning techniques during VLM training, utilizing similar manipulated data as hard negatives. However, despite the fine-tuning process, false positive rewards remain unavoidable.
@@
Figure 11: Example of training data for the Montezuma environment.
We used the threshold $\hat{q}$ introduced in Section 6.1 to make binary classification on the testing pairs to evaluate the performance of the fine-tuned VLM-based reward models. We found that VLM models had difficulty achieving high accuracy on Minigrid environment, which is likely due to the too abstract and cartoonish nature of the environment, causing the VLMs to struggle to learn the visual-textual correspondence. We also found that X-CLIP did not perform better than CLIP in our experiments. We hypothesize that the cartoonish nature of the testing environments may have caused the X-CLIP model to struggle to learn the visual-textual correspondence. Thus, we used CLIP as the backbone model throughout our following experiments. The performance of the fine-tuned VLM-based reward models is shown in Table 3. Even when the precision score reaches 0.98, indicating that only 2% of the rewards are false positives in the validation set, the agent can still significantly underperform in the testing environments. The core issue is that in out-of-distribution (O.O.D.) testing environments, false positive rewards are prevalent and inevitable. Therefore, it is crucial to design a reward function that is robust to reward noise.
@@
A.4.4 Hyperparameters for Instruction-Following RL Agents
In the experiments, all methods are implemented based on PPO with same model architecture. The Minigrid and Crafter environments use the same training hyperparameters as the Achievement Distillation paper (Moon et al., 2023). For Montezuma’s Revenge, we found that the performance of the agent was sensitive to the gamma and GAE lambda parameters. To improve the performance of agents in Montezuma’s Revenge, we took two additional steps: (1) normalizing the observation inputs when computing the rewards, and (2) not normalizing the advantage during the GAE calculation. The hyperparameters are shown in the following tables.
Table 4: Model Parameters

### v5

@@ Appendix D Convergence time on a sparse-reward landscape
Proof.
Let $\vec{X_{1}}$, $\vec{X_{2}}$, … $\vec{X_{T}}$ be IID random unit vectors uniformly distributed on a (d1)(d-1)-dimensional sphere $\mathcal{S}^{d-1}\subset\mathbb{R}^{d}$, where $\vec{X_{i}}=(X_{i1},...,X_{id})$ and $|\vec{X_{i}}|^{2}=\sum_{j=1}^{d}X^{2}_{ij}=1$. Let $\vec{S_{T}}:=\sum_{i=1}^{T}\vec{X_{i}}$. By a $n$-dimensional cosine rule we have
@@
[…]
@@
See 4.2
Proof.
The expected time taken for each of the sub-tasks then scales with $d_{i}^{2}$ respectively (Lemma D.1), and we have
@@
because $d_{i}>0\forall i$. We can also work out that the upper bound for this improvement is a factor of n1n-1 by invoking the Cauchy–Schwarz inequality $\left(\sum_{i=1}^{n}u_{i}^{2}\right)\left(\sum_{i=1}^{n}v_{i}^{2}\right)\geq\left(\sum_{i=1}^{n}u_{i}v_{i}\right)^{2}$:
@@
and the equality sign holds (indicating maximal improvement) when $d_{1}=d_{2}=...=d_{n-1}$. The intuitive interpretation is that the divide-and-conquer approach is the most effective when the task is divided evenly into subtasks. ∎
Appendix E VLM-based reward model is a heuristic function
See 4.4
Proof.
Consider a sparse reward environment where the external reward is defined as:
$r^{e}(s_{t})=\begin{cases}1&\text{if }s_{t}\in\mathcal{S}_{G},\\ 0&\text{otherwise},\end{cases}$
with discount factor $\gamma\in(0,1)$. For an optimal policy $\pi^{*}$, the value function at state $s_{t}$ is:
$V^{*}(s_{t})=\gamma^{\widetilde{T}-t},$
where $\widetilde{T}$ is the time to reach a goal state $s_{g}\in\mathcal{S}_{G}$ along the shortest path from $s_{0}$, and $\widetilde{T}-t$ is the remaining steps from $s_{t}$. This follows because the agent receives a single reward of 1 at $s_{g}$, discounted back to $s_{t}$.
In Algorithm 1, the reward $r^{v}(\tau_{t},l_{m(t)})\in[0,1]$ measures alignment between the agent’s transition $(\tau_{t-1},s_{t},a_{t})$ and the expert instruction $l_{m(t)}$ for sub-task $m(t)$. By design, $r^{v}=1$ when the transition fully matches $l_{m(t)}$, and $r^{v}<1$ otherwise, implying partial alignment.
We assume expert instructions $\{l_{1},\dots,l_{n}\}$ define a trajectory $\tau^{*}$ that reaches $\mathcal{S}_{G}$ optimally or near-optimally. If the agent fully follows $l_{m(t)}$ at each step, its path aligns with $\tau^{*}$, and $\widetilde{T}$ is minimized (say, $\widetilde{T}^{*}$). When the agent perfectly follows the instruction at each step along $\tau^{*}$, the VLM reward is maximized, i.e., $r^{v}(\tau_{t},l_{m(t)})=1$. Crucially, moving along the optimal trajectory $\tau^{*}$ leads to the optimal value function $V^{*}(s_{t})=\gamma^{\widetilde{T}^{*}-t}$.
Now, if the agent deviates from $l_{m(t)}$, it takes a suboptimal action, increasing the expected path length to $\mathcal{S}_{G}$. Let $\widetilde{T}^{\prime}>\widetilde{T}^{*}$ denote this detour. Then, $V^{*}(s_{t})=\gamma^{\widetilde{T}^{\prime}-t}<\gamma^{\widetilde{T}^{*}-t}$. Since $r^{v}$ decreases with misalignment (e.g., from 1 to some $0\leq r<1$), it exhibits a monotonic relationship with the value function: higher alignment (larger $r^{v}$) corresponds to movement towards a shorter path and thus a larger $V^{*}(s_{t})$, and vice versa.
Thus, VLM-based rewards, under Algorithm 1, function as a heuristic $h(s_{t})$ that approximates the relative ordering of states in terms of their optimal value $V^{*}(s)$. ∎
Appendix F False Positive and the Violation of Pessimistic Property of Heuristics
See 4.9
@@ Appendix F False Positive and the Violation of Pessimistic Property of Heuristics
    Express the Bellman equation for $h$: $(\mathcal{B}h)(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)h(s^{\prime})$
@@
Appendix G False Positive gives unbounded bias in the optimality gap
In this work, Section 4, we analyze the impact of false positive rewards on the convergence speed of policy learning based on the optimality gap analysis by Cheng et al. (2021). Here, we provide a detailed proof of the impact of false positive rewards on the convergence speed of policy learning.
We have defined the formal definition of false positive and false negative from both instruction-following (IF) and heuristic perspectives in Section 4.
See 4.10 We analyze the convergence speed of policy learning by examining the optimality gap, which is defined as the difference between the optimal value of the initial state $s_{0}$, $V^{*}(s_{0})$, and the value of the initial state under an arbitrary policy $\pi$, $V^{\pi}(s_{0})$. Specifically, we focus on deriving an upper bound for this optimality gap. The key intuition is that a smaller upper bound implies faster convergence to the optimal policy, as fewer iterations of policy updates will be required to reach the optimum.
@@ Appendix G False Positive gives unbounded bias in the optimality gap
Theorem G.1 (Optimality Gap Decomposition (Cheng et al. 2021)).
@@
We will not provide a proof for Theorem G.1 in this paper. Please refer to (Cheng et al. 2021) for details. The theorem demonstrates that the performance gap can be elegantly decomposed into two components: a regret term and a bias term such that:
@@
Cheng et al. (2021) have further derived the upper bound for the bias term, which we present here as a lemma. We omit the proof for brevity; for detailed proof, please refer to (Cheng et al. 2021).
Lemma G.2 (Upper Bound of the Bias Term).
@@
    To begin with, we need to use another lemma from (Cheng et al. 2021), shown as follows:
    Lemma G.4 (Bellman equation of reshaped MDP (Cheng et al. 2021)).
@@
    For brevity, we do not include the proof for Lemma G.4.
@@
    Then, according to Lemma G.4, we can get
@@
    Recall that the use of VLM-based reward model transform the original MDP into a shaped MDP $\widetilde{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},\tilde{r},\gamma\rangle$, where the reward function becomes $\tilde{r}=r^{e}(s,a)+r^{v}(\tau_{t},l_{m(t)})$. Also, it is important to distinguish the Bellman backup equation under the two different MDPs: the original $\mathcal{M}$ and the reshaped $\widetilde{\mathcal{M}}$. By definition, $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$. In contrast, $(\widetilde{\mathcal{B}}h)(s,a)=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$. Nevertheless, the two Bellman backup equations possess a remarkable property – they are essentially equivalent to each other:
    $(\widetilde{\mathcal{B}}h)(s,a)$ $=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
    $=(\mathcal{B}h)(s,a)$ (34)
    Let $\pi$ denote the greedy policy of $\arg\max_{a}(\mathcal{B}h)(s,a)$ and then trace back to Equation 34, we have $(\widetilde{\mathcal{B}}h)(s,a)=(\mathcal{B}h)(s,a)$, that means
@@
    $\geq h(s_{0})\quad\text{[direct result of $h$ being pessimistic]}$ (37)
    Generalization to Any State: The lemma from (Cheng et al. 2021) (Lemma G.4) is stated in terms of any starting state $s_{0}$. Therefore, we can replace $s_{0}$ with any state $s\in\mathcal{S}$ in our analysis. Thus we get $\forall s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
    Lemma G.3 implies that when $h$ is pessimistic, $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\leq 0$ because $\widetilde{V}^{*}(s)\geq h(s)$ for all $s$. Therefore, the bias error is only bounded by $B_{1}$ when $h$ is pessimistic, i.e., $\mathrm{Bias}(h,\beta,\pi)\leq B_{1}$.
@@
    We show briefly that, unlike the underestimation error, the overestimation error does not have a closed-form upper bound expression. The difficulty, as pointed out by Cheng et al. (2021), originates from the trajectory-dependence on $\mathbb{E}_{\rho^{\pi}}[\cdot]$ as the $l_{\infty}$ approximation error here can be difficult to control in large state spaces. It is possible that, upon picking up falsely high $h$ states, $\rho^{\pi}$ gets further distorted away from $\rho^{\pi^{*}}$ and accumulates more falsely high $h$ states. In other words, this trajectory dependence makes $B_{2}$ prone to a feedback loop of accumulating overestimation errors and resulting in a much larger upper bound for the bias.
G.1 Convergence Guarantee of HuRL Compared With Potential-Based Reward Shaping
Previous work on potential-based reward shaping (PBRS) has established that learning a policy on a reshaped MDP can converge to the optimum of the original MDP, as proved by Ng, Harada, and Russell (1999). They proved that the optimal $Q$ value of the reshaped MDP is equivalent to the optimal $Q$ value of the original MDP minus a state-dependent function. Consequently, for an optimal policy defined as $\pi^{*}=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a)$, this additional state-dependent value does not alter the policy. Thus, the optimal policy of the original MDP $\mathcal{M}$ is also the optimal policy for the reshaped $\widetilde{\mathcal{M}}$. However, the theoretical convergence for the reshaped MDP within the framework of Heuristic-Guided Reinforcement Learning (HuRL) has not been proven with the same rigor.
The authors of HuRL made a trick which involves manipulating the coefficient $\beta$, which scales the heuristic-based auxiliary rewards. If $\beta$ increases gradually from 0 to 1 over the training process, the agent effectively transitions from interacting with the modified MDP back to the original MDP. This method ensures that, at the end of training, the agent is exactly solving the original problem. Moreover, according to the Blackwell optimal property (Chen, Kochenderfer, and Spaan 2018), convergence can occur before $\beta$ reaches 1, thereby allowing the trained policy to maintain optimality in the original MDP under HuRL.
@@ G.1 Convergence Guarantee of HuRL Compared With Potential-Based Reward Shaping
Appendix H Additional Experimental Setup Details
H.1 VLM-based Reward Model Procedure
@@
We provide a brief overview of the VLM-based reward model procedure. For a detailed procedure, please refer to Appendix C. The VLM-based reward model will have a pointer to the sequence of the instruction sentence, starting at the first sentence. For original models Pixl2R and ELLM-, we follow the setting in their original work where for each instruction sentence (the full instruction essay will be split into multiple sentences and treat each sentence as atomic instruction $l_{k}$), the reward model will have a maximum cap of rewards (2.0) it can assign to the agent in one episode. When the cap is reached, the reward model will move its pointer to the next instruction sentence. This cap value was selected as a hyperparameter in prior work, intended to serve as a heuristic threshold indicating instruction completion. However, we note that this approach does not guarantee semantic completion, as high cumulative similarity scores can still occur despite incomplete or incorrect execution. For the BiMI reward model, the reward model will move its pointer to the next instruction sentence when the binary signal is triggered.
In terms of the compute resources, we mainly use one NVIDIA 3090 GPU (24GB VRAM) for training and running both the VLM-based reward model and the RL policy model.
H.2 Finetuning VLM-based Reward Models
@@
In contrast to previous work where they rely on hand-crafted oracle multimodal reward models (e.g., Du et al. (2023)), we use actual pretrained VLMs to generate reward signals. 2 VLM backbone models are used in our experiments: 1) CLIP (Radford et al. 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al. 2022), pretrained by video-text pairs. In particular, Pixl2R uses CLIP because it only uses the single latest frame as input. In contrast, ELLM- takes a slice of trajectory (i.e., multiple frames) as input, and thus uses either X-CLIP or CLIP with additional RNN encoder as the reward model.
@@
Due to the cartoonish and abstract visuals of the testing environments, we further fine-tune the VLMs to adapt to this new visual domain. This fine-tuning is based on the contrastive learning method. To collect data, we use well-trained expert agents from (Moon et al. 2023) to generate expert trajectories for the Crafter environments and annotate them with instructions using internal information from the game engine. For Minigrid environments, we use classical search-based planning robots to generate expert trajectories and annotate them with the corresponding task instructions. For Montezuma’s Revenge, we manually annotate the expert trajectories.
For Minigrid and Crafter, we collected approximately 80,000 training pairs, each consisting of a natural language instruction and a corresponding visual observation (i.e. a short trajectory snippet). These annotations were generated using internal game state information with corresponding instruction templates. For Montezuma’s Revenge, we manually created around 300 high-quality instruction-observation pairs based on expert demonstrations. These training data are of high quality, as we have made every effort to avoid false positive rewards due to poor training data quality. However, despite the fine-tuning process, false positive rewards remain unavoidable. Figure 11 presents an example of instruction data employed for VLM finetuning in the Montezuma’s Revenge environment.
@@
Figure 11: Example of expert instruction data for VLM finetuning under in Montezuma’s Revenge environment.
We determine the threshold $\hat{q}$ using conformal prediction to convert continuous cosine similarity scores into binary labels for evaluating the VLM’s ability to recognize correct instruction-observation pairs. This allows us to compute standard classification metrics such as precision, recall, and F1-score—providing a preliminary assessment of VLM performance independent of downstream RL interaction. The performance of the fine-tuned VLM-based reward models is shown in Table 4. However, this high precision does not guarantee robust agent performance. As we demonstrate later, agents guided by these models significantly underperform in out-of-distribution (O.O.D.) testing environments, where false positive rewards become prevalent due to distributional shifts, revealing a critical limitation in generalization.
@@
H.3 Hyperparameters for VLM Reward Model + RL Agents
In the experiments, all methods are implemented based on PPO with same model architecture. The Minigrid and Crafter environments use the same training hyperparameters as the Achievement Distillation paper (Moon et al. 2023). For Montezuma’s Revenge, we found that the performance of the agent was sensitive to the gamma and GAE lambda parameters. To improve the performance of agents in Montezuma’s Revenge, we took two additional steps: (1) normalizing the observation inputs when computing the rewards, and (2) not normalizing the advantage during the GAE calculation. The hyperparameters are shown in Table 5, Table 6 and Table 7.
Table 5: Agent Policy Model Parameters